### v1

# Resilient by Design – Active Inference for Distributed Continuum Intelligence

Abstract
Failures are the norm in highly complex and heterogeneous devices spanning the distributed computing continuum (DCC), from resource-constrained IoT and edge nodes to high-performance computing systems. Ensuring reliability and global consistency across these layers remains a major challenge, especially for AI-driven workloads requiring real-time, adaptive coordination. This paper introduces a Probabilistic Active Inference Resilience Agent (PAIR-Agent) to achieve resilience in DCC systems. PAIR-Agent performs three core operations: (i) constructing a causal fault graph from device logs, (ii) identifying faults while managing certainties and uncertainties using Markov blankets and the free-energy principle, and (iii) autonomously healing issues through active inference. Through continuous monitoring and adaptive reconfiguration, the agent maintains service continuity and stability under diverse failure conditions. Theoretical validations confirm the reliability and effectiveness of the proposed framework.
I Introduction
No one can deny that computing infrastructure has transformed dramatically with Generative AI, driving unprecedented demand for high-performance resources spanning sensor nodes, edge devices, network elements, and server clusters. In response, the distributed compute continuum (DCC) paradigm has emerged, orchestrating complex tasks across the full device continuum, allocating smaller, latency-sensitive tasks locally to reduce energy use and preserve privacy, while offloading resource-intensive tasks to high-performance servers [1, 2]. Such dynamic task distribution optimizes resource use, balances energy demand, supports scalable operations. As the number and diversity of devices increase, data volumes surge, and workloads become more dynamic, system complexity concurrently grows, intensifying failure risks [3]. In general, task processing failures in the DCC occur due to factors such as user-initiated aborts, incorrect inputs, insufficient allocated resources, execution errors, node crashes, or communication failures; In DCC, additional failures include the unpredictability brought by device heterogeneity, mobile and battery-dependent nodes, dynamic network and resource shifts, and large-scale consistency challenges associated with hundreds or thousands of edge devices.
Resilience mechanisms are crucial for modern computing systems, since failures are the norm in these infrastructures. The term resilience is chosen over mere fault tolerance in DCC because these systems must not only tolerate discrete faults but also effectively manage and adapt to a wide range of inherent uncertainties [4, 5]. In DCC, resilience reflects the system’s capability to sustain continuous operation through proactive adaptation, localized recovery, and graceful degradation across diverse cloud, fog, edge, and IoT layers, where unpredictable workload fluctuations, device mobility, connectivity disruptions, and energy variability are common [6, 7]. To achieve these, this paper designs a novel probabilistic active inference resilience agent (PAIR-Agent), with the following contributions.
  • PAIR-Agent builds a probabilistic Causal Fault Graph (CFG) using Bayesian Network Structure Learning (BNSL) through various log data from the devices.
  • Next, it uses the Markov blanket to identify potential faults from CFGs. Further, PAIR-Agent applies the free-energy principle for uncertainty-aware fault inference.
  • Finally, PAIR-Agent autonomously selects corrective actions minimizing expected free energy and enabling self-healing through active inference.
II System Model
The DCC ($\mathcal{S}$) are composed of $n$ devices including one or $c$ cloud or data centers $\mathcal{C}=\{C_{1},C_{2},...,C_{c}\}$, a set of $f$ Fog nodes $\mathcal{F}=\{F_{1},F_{2},...,F_{f}\}$, a set of e|e|Edge nodes $\mathcal{E}=\{E_{1},E_{2},...,E_{e}\}$, $m$ mobile devices $\mathcal{M}=\{M_{1},M_{2},...,M_{m}\}$, a set of $\iota$ IoT devices $\Gamma=\{I_{1},I_{2},...,I_{\iota}\}$, and a set of $\varsigma$ sensor nodes (SNs) $\Psi=\{\psi_{1},\psi_{2},...,\psi_{\varsigma}\}$. We assume that $\Psi$ are data producers and they are not capable of computing any data. Therefore, they transmit their data to the nearest base station, i.e., $\Gamma$, $\mathcal{M}$, or $\mathcal{E}$ using wireless communication through Bluetooth Low Energy, Zigbee, Cellular, or Wi-Fi channels [8].
A set of AI-related computational tasks is defined as $\mathcal{T}=\{T_{1},T_{2},...,T_{t}\}$, where each task $T_{i}=\langle\omega_{i},\chi_{i},\phi_{i},\delta_{i}\rangle$ denotes workload $\omega_{i}$, input data dependencies $\chi_{i}$, a computational mapping or model function $\phi_{i}:\chi_{i}\mapsto y_{i}$, and an execution deadline $\delta_{i}$. The DCC graph $\mathcal{S}=(\mathcal{V},\mathcal{R})$ consists of nodes $\mathcal{V}=\{\mathcal{C}\cup\mathcal{F}\cup\mathcal{E}\cup\mathcal{M}\cup\Gamma\cup\Psi\}$ and communication relationships $\mathcal{R}\subseteq\mathcal{V}\times\mathcal{V}$, where each node $v_{j}\in\mathcal{V}$ executes a subset of tasks $\mathcal{T}(v_{j})$ such that $\bigcup_{v_{j}\in\mathcal{V}}\mathcal{T}(v_{j})=\mathcal{T}$. Each task execution $T_{i}(v_{j})$ follows a three-phase structure $\langle\alpha_{i},\beta_{i},\gamma_{i}\rangle$, representing task initiation, computation, and completion or checkpointing, respectively. In general, these complex AI workflows such as training are composed of hierarchically nested subtasks, e.g., $T_{\text{train}}=T_{\text{data-load}}\triangleright T_{\text{forward}}\triangleright T_{\text{backward}}\triangleright T_{\text{update}}$, where each subtask can complete and save progress independently to support partial recovery. Each task establishes a stable checkpoint $\zeta_{i}=\langle s_{i},D_{i},\theta_{i}\rangle$, recording the intermediate state $s_{i}$, processed dataset segment $D_{i}$, and model parameters $\theta_{i}$, enabling restoration and continuation upon failure.
III Resilience in Action
The proposed PAIR-Agent iteratively performs the following steps (as depicted in Figure 1): (a) identifying changes in the system logs and collecting data to generate a CFG; (b) performing fault inference by determining certain and uncertain faults using the Markov blanket and free energy principle[9]; and (c) healing the faults through active inference to achieve resilience, as discussed in more detail in the subsequent subsections.
Figure 1: The proposed PAIR-Agent for resilience in the DCC, collects and normalizes logs to form a CFG, infers certain and uncertain faults via Markov blanket and FEP analysis, and healing through active inference to restore system stability.
III-A Collect and Process Logs
Initially, the PAIR-Agent coordinates distributed log acquisition across the $\mathcal{S}$ using remote procedure calls (RPCs) to query each device $(v_{i}\in\mathcal{S})$ for its most recent checkpoint $(\zeta_{i}^{t}=\langle s_{i}^{t},D_{i}^{t},\theta_{i}^{t}\rangle)$, which encapsulates the local state, processed dataset segment, and learned parameters at analysis round $t$. These checkpoints act as temporal anchors that partition system observations into discrete analysis intervals, allowing the agent to process only the log entries generated after $\zeta_{i}^{t}$. This design ensures synchronization across heterogeneous nodes, prevents redundant ingestion of log segments, tracks round-based progress, and maintains a globally consistent causal timeline of system activity [10]. Each new round $(t+1)$ therefore, begins with a synchronized collection of incremental log updates $\Delta L_{i}^{t+1}$, forming a time-evolving dataset that supports continual inference.
[…]
To uncover causal dependencies among probabilistic fault variables, PAIR-Agent constructs a time-evolving CFG $\mathcal{G}_{\text{fault}}^{t+1}=(\mathcal{V}_{f}^{t+1},\mathcal{E}_{f}^{t+1})$ using a BNSL. Here, each node $v_{i}\in\mathcal{V}_{f}^{t+1}$ represents a stochastic variable corresponding to a specific fault or contextual feature extracted from the normalized feature matrix $X^{t+1}$, and each directed edge $e_{ij}\in\mathcal{E}_{f}^{t+1}$ encodes a conditional dependency $P^{t+1}(v_{j}\mid\text{Pa}(v_{j}))$, where $\text{Pa}(v_{j})$ denotes the parent set of $v_{j}$. This structure is inferred incrementally at each checkpoint interval $(t,t+1]$, allowing PAIR-Agent to integrate newly observed evidence $\Delta\mathcal{D}^{t+1}$ while retaining prior knowledge from earlier analysis rounds. This is achieved through Hill-Climbing with Bayesian Dirichlet equivalence scoring [11], combined with probabilistic regularization to handle noise and partial observability in DCC data. In scenarios with strong nonlinearity or non-stationary correlations, PAIR-Agent employs variational Bayesian learners to adaptively refine causal relations across checkpoints. Each updated Bayesian network $\mathcal{G}_{\text{fault}}^{t+1}$ represents the posterior belief over causal structure given new observations, expressed as $P^{t+1}(\mathcal{G}_{\text{fault}}\mid\mathcal{D}^{t+1})\propto P(\mathcal{D}^{t+1}\mid\mathcal{G}_{\text{fault}})\,P^{t}(\mathcal{G}_{\text{fault}}),$ where $P^{t}(\mathcal{G}_{\text{fault}})$ encodes the prior structural belief accumulated from previous rounds. The resulting sequence of CFGs $\{\mathcal{G}_{\text{fault}}^{1},\mathcal{G}_{\text{fault}}^{2},\dots,\mathcal{G}_{\text{fault}}^{t+1}\}$ provides a temporally consistent and probabilistically grounded representation of causal fault propagation.
III-B Fault Inference
Building upon the Bayesian CFG $\mathcal{G}_{\text{fault}}^{t+1}$, PAIR-Agent performs probabilistic fault inference by exploiting the local Markov properties of the network. For each fault variable $f_{i}\in\mathcal{V}_{f}^{t+1}$, the agent defines its Markov blanket $\mathcal{M}(f_{i})=\text{Pa}(f_{i})\cup\text{Ch}(f_{i})\cup\text{Pa}(\text{Ch}(f_{i}))$, where $\text{Pa}(f_{i})$ and $\text{Ch}(f_{i})$ denote the parent and child sets of $f_{i}$, respectively. This blanket represents the minimal sufficient set of variables that render $f_{i}$ conditionally independent from the rest of the network. By monitoring the probabilistic interactions within $\mathcal{M}(f_{i})$, the agent computes the posterior likelihood of fault activation or propagation, i.e., $P^{t+1}(f_{i}\mid\mathcal{M}(f_{i}))\propto P^{t+1}(f_{i}\mid\text{Pa}(f_{i}))\prod_{c_{j}\in\text{Ch}(f_{i})}P^{t+1}(c_{j}\mid\text{Pa}(c_{j}))$.
At each checkpoint interval $(t,t+1]$, the agent continuously updates these posteriors using newly collected evidence $\Delta\mathcal{D}^{t+1}$, distinguishing between certain and uncertain faults. A fault $f_{i}$ is classified as certain when its posterior confidence exceeds a threshold $\tau_{c}$, i.e., $P^{t+1}(f_{i}\mid\mathcal{M}(f_{i}))\geq\tau_{c}$, indicating high evidence consistency across nodes. Conversely, when the probability lies within the uncertainty band $\tau_{u},\tau_{c})$, it is treated as an uncertain fault, requiring further active inference and data acquisition. To refine fault-type attribution, PAIR-Agent further decomposes the inference space into hardware-related and software-related variables i.e., $\mathcal{V}_{f}^{t+1}=\mathcal{V}_{\text{HW}}^{t+1}\cup\mathcal{V}_{\text{SW}}^{t+1},$ where $\mathcal{V}_{\text{HW}}^{t+1}$ includes factors such as temperature [12], power fluctuation, or connectivity degradation (this can be both hardware or software based on type), and $\mathcal{V}_{\text{SW}}^{t+1}$ captures task failures, model divergence, or resource contention anomalies. Through conditional marginalization, the system derives joint distributions i.e., $P^{t+1}(f_{i}\mid\mathcal{V}_{\text{HW}}^{t+1})\quad\text{and}\quad P^{t+1}(f_{i}\mid\mathcal{V}_{\text{SW}}^{t+1}),$ allowing discrimination between underlying fault origins and supporting targeted recovery strategies. $\mathcal{V}_{\text{HW}}^{t+1}$ are managed through both human intervention and automated software responses, depends on the kind of fault.
Physical faults, like disconnected cables or damage of the device, require manual repair, while software-manageable issues, such as dropped Wi-Fi, can be mitigated remotely via alternative networks or redundancy. Automated measures, like reducing load or adjusting cooling, handle overheating issues. Further, PAIR-Agent integrates free-energy minimization with the Bayesian CFG and Markov blanket structure to quantify fault likelihoods while accounting for interdependencies and hidden uncertainties. Under uncertainties, PAIR-Agent utilizes the free-energy principle (FEP) to continuously align its internal model with observed system behavior. At each checkpoint $t+1$, the variational free energy is defined as $F^{t+1}=\mathbb{E}_{q^{t+1}(f_{i})}\Big[\ln q^{t+1}(f_{i})-\ln P^{t+1}(f_{i},\mathbf{x}_{i}^{t+1})\Big],$ where $q^{t+1}(f_{i})$ approximates the posterior probability of fault $f_{i}$ given the contextual features $\mathbf{x}_{i}^{t+1}$. Minimizing Ft+1 updates these beliefs, effectively refining the agent’s estimates of both certain and uncertain faults. With FEP, PAIR-Agent can quantify fault likelihoods while accounting for interdependencies and hidden uncertainties.
III-C Active Inference for Healing
Once fault inference identifies and classifies faults with associated uncertainties, PAIR-Agent employs active inference [13] to autonomously plan and execute corrective actions. The agent evaluates a set of possible actions $a_{t}\in\mathcal{A}$ and selects the one that minimizes the expected divergence between predicted and desired system states, formalized as $a_{t}^{*}=\arg\min_{a_{t}}G(a_{t}),\quad G(a_{t})=\mathbb{E}_{Q(s_{t+1}\mid a_{t})}\Big[D_{\mathrm{KL}}\!\left(Q(s_{t+1})\parallel P(s_{t+1})\right)-\ln P(o_{t+1}\mid s_{t+1})\Big],$ where $Q(s_{t+1}\mid a_{t})$ denotes the predicted distribution of future system states under action $a_{t}$, $P(s_{t+1})$ encodes the desired operational state, and $P(o_{t+1}\mid s_{t+1})$ represents the likelihood of observing evidence consistent with a healthy configuration. By minimizing $G(a_{t})$, PAIR-Agent aligns system behavior with target operational objectives while accounting for uncertainties across both hardware and software variable sets $\mathcal{V}_{\text{HW}}^{t+1}$ and $\mathcal{V}_{\text{SW}}^{t+1}$. Each candidate action is further assessed for its expected impact on the inferred fault probabilities. The anticipated reduction in fault likelihood is modeled as $\hat{P}(f_{i}\mid a_{t})=P(f_{i})\cdot(1-w_{i}(a_{t})),$ where $w_{i}(a_{t})$ quantifies the effectiveness of action $a_{t}$ in mitigating fault $f_{i}$. The optimal corrective action minimizes the expected residual risk across all active faults: $a_{t}^{*}=\arg\min_{a_{t}\in\mathcal{A}}\sum_{i}\hat{P}(f_{i}\mid a_{t}),$ ensuring effective and energy-efficient mitigation of disruptions. After executing the selected action, PAIR-Agent observes the resulting system state, updates posterior fault probabilities $P^{t+1}(f_{i}\mid\mathbf{x}_{i}^{t+1})$ with new evidence, and advances the checkpoint $\zeta^{t+1}$ to record the updated configuration and belief state. This closed-loop cycle of perception–inference–action–update enables continual adaptation and resilience across analysis rounds. During the healing process, PAIR-Agent not only redistributes tasks but also performs device-level recovery through automated reinitialization, firmware reloading, or adaptive configuration tuning. If such self-healing measures fail to restore stability, the agent isolates the faulty node and applies corrective actions at the continuum level $\mathcal{S}$, including workload redistribution across fog or cloud layers, rerouting through redundant network paths, managing dynamic load or thermal management, and escalation to human operators for physical interventions.
IV Theoretical Results and Discussions
This section presents the theoretical foundations demonstrating that the proposed PAIR-Agent framework achieves resilient operation in the DCC through probabilistic reasoning, causal inference, and active self-healing. Three main results establish its correctness and convergence properties.
Result 1 (Convergence of Belief Updating). Let $P^{t}(f_{i})$ denote the prior fault belief at checkpoint $t$ and $Q^{t+1}(f_{i})$ its posterior after processing new evidence $\Delta\mathcal{D}^{t+1}$. Under the Bayesian update rule $Q^{t+1}(f_{i})\propto P(\Delta\mathcal{D}^{t+1}\mid f_{i})P^{t}(f_{i}),$ the Kullback–Leibler divergence between successive posterior estimates satisfies $D_{\mathrm{KL}}\!\left(Q^{t+1}(f_{i})\parallel Q^{t}(f_{i})\right)\to 0\quad\text{as }|\Delta\mathcal{D}^{t+1}|\to 0.$ Hence, as log updates become incremental, the posterior distribution converges almost surely to the true fault probability $P^{*}(f_{i})$. This ensures consistent belief refinement across checkpoints, guaranteeing stable and bounded fault estimation in the presence of streaming data.
Result 2 (Optimal Action Selection under Expected Free Energy Minimization). Let $G(a_{t})$ denote the expected free energy associated with action $a_{t}\in\mathcal{A}$: $G(a_{t})=\mathbb{E}_{Q(s_{t+1}\mid a_{t})}\Big[D_{\mathrm{KL}}\!\left(Q(s_{t+1})\parallel P(s_{t+1})\right)-\ln P(o_{t+1}\mid s_{t+1})\Big].$ Assuming $G(a_{t})$ is convex in the neighborhood of the optimal policy and both $Q(s_{t+1}\mid a_{t})$ and $P(s_{t+1})$ are Lipschitz continuous, the selected action $a_{t}^{*}=\arg\min_{a_{t}}G(a_{t})$ minimizes the upper bound of predictive uncertainty while maximizing the evidence of expected observations. Thus, PAIR-Agent’s policy converges to an equilibrium minimizing both epistemic (uncertainty-driven) and pragmatic (goal-driven) free energy, ensuring that corrective actions asymptotically restore the DCC to its desired operational manifold.
Result 3 (System-Level Resilience via Active Healing Loop). Consider the system state dynamics $s_{t+1}=f(s_{t},a_{t},\xi_{t})$, where $\xi_{t}$ represents random disturbances or failures. Define resilience as the probability of maintaining bounded deviation from the nominal state, i.e., $\mathbb{P}(\|s_{t+1}-s^{*}\|\leq\epsilon)\geq 1-\delta,$ for tolerance $\epsilon$ and confidence $1-\delta$. Under PAIR-Agent’s closed-loop policy that minimizes both expected free energy $G(a_{t})$ and residual fault probability $\sum_{i}\hat{P}(f_{i}\mid a_{t})$, the composite Lyapunov function $V_{t}=F^{t}+\lambda\sum_{i}\hat{P}(f_{i})$ is non-increasing for all $t>0$, the mean-square boundedness of $s_{t}$. Therefore, the DCC remains resilient against stochastic faults, achieving self-stabilization through the proposed PAIR-Agent.
V Conclusions
This paper introduced the PAIR-Agent as a unified framework for achieving resilience in DCC systems. The PAIR-Agent agent performs: (i) constructing a CFG through device log analysis, (ii) identifying faults while managing certainties and uncertainties using Markov blankets and the free-energy principle, and (iii) autonomously healing issues via active inference. Through its closed-loop cycle of observation, inference, and action, PAIR-Agent ensures adaptive stability, self-healing capability, and sustained operational continuity, marking a significant step toward fully autonomous and resilient DCC systems. Theoretical results confirm its reliability, and future work will extend this research with detailed and distributed implementation, empirical validation, and evaluation under real-time environments.